OpenStudy (anonymous):

let cscx = -5/3 and cosx < 0, find tan x/2

OpenStudy (e.mccormick):

So what part of this are you having trouble with?

OpenStudy (anonymous):

all of it. haha. just need help to solve it, please.

OpenStudy (e.mccormick):

You have \(\cos x < 0\) which means it is negative. What does that tell you about the possible answers?

OpenStudy (anonymous):

negative.

OpenStudy (e.mccormick):

Well, it is for tan in the end... so they might or might not. But what quadrent(s) would the answers be in?

OpenStudy (anonymous):

quadrant 3.

OpenStudy (e.mccormick):

Is cos x negative anywhere other then Q3?

OpenStudy (anonymous):

i have two more that relate to the same question but the two are find secx/2 and find cotx

OpenStudy (e.mccormick):

Or are you saying that absed off of both the cscx = -5/3 and cosx < 0 it is Q3. I was tak9ng it from the start because you said you had trouble with all of it.

OpenStudy (e.mccormick):

missed the i there. lol. Hit 9.

OpenStudy (anonymous):

try working it out, so i can compare it to the answer i got.

OpenStudy (e.mccormick):

These are basically about building a triangle on the unit circle. The sides of the triangle show the relationship of the trig ratios. Once you know which quad, then it is very easy to sketch and then pull all the answers from the sketch.

OpenStudy (e.mccormick):

Well, what answers did you get?

OpenStudy (anonymous):

hold on.

OpenStudy (e.mccormick):



OpenStudy (e.mccormick):

As you can see from that, the x value will be -4.

OpenStudy (e.mccormick):

Things started reloading and lagging, so I never got to see what you had for an answer. I had this all typed up! Well, here it is in case you still need that checked. You have \(\cos x < 0\) which means it is negative. What does that tell you about the possible answers? sin csc cos sec \(y=-3\) \(r=5\) \(x=\sqrt{5^2-(-3)^2}\implies\sqrt{16}\implies\pm4\) and we know it is negative. \[\begin{matrix} \sin\theta=\frac{y}{r}&\csc\theta=\frac{r}{y}\\ \cos\theta=\frac{x}{r}&\sec\theta=\frac{r}{x}\\ \tan\theta=\frac{y}{x}&\cot\theta=\frac{x}{y} \end{matrix}\\ \therefore \\ \begin{matrix} \sin\theta=\frac{-3}{5}&\csc\theta=\frac{5}{-3}\\ \cos\theta=\frac{-4}{5}&\sec\theta=\frac{5}{-4}\\ \tan\theta=\frac{-3}{-4}&\cot\theta=\frac{-4}{-3} \end{matrix}\]Now, some of those have two negatives, which would cancel and have an end result of a positive.